Three dice are thrown. Find the probability that the first one shows a five, and the last two each show even numbers.
Answer by Edwin McCravy(20065)

There are 6 numbers on a die. They are 1,2,3,4,5,6.
The probability of a die showing a 5 is 1/6.
The even numbers on a die are 2,4, and 6
The probability of a die showing an even number is 3/6 which reduces to 1/2.

Answer: P(5EE) = (1/6)(1/2)(1/2) = 1/24

Edwin